# Long-term stable nonlinear evolutions of ultracompact black-hole mimickers

### v1

@@ Appendix E Ray tracing
The ray-traced images are obtained using FOORT, a Flexible Object-Oriented Ray Tracer that is publicly available on GitHub Mayerson (2024); Mayerson et al. . The geodesic equations are integrated using the Verlet scheme in order to produce images of a variety of spacetimes. It is highly flexible and easily accommodates new analytical or numerical spacetimes, such as a boson star. It employs adaptive mesh refinement. The images in the main text are produced assuming an optically thin equatorial disk with emission profile given by
$I_{0}(r)=\,$ $\left((r-\rho)^{2}+\Delta^{2}\right)^{-1/2}$
@@
×$\times\exp\left(-\frac{1}{2}\left[\kappa+\operatorname{arcsinh}\left(\frac{r-\rho}{\Delta}\right)\right]^{2}\right)\,,$ (30)
@@
a simple analytical model proposed in Ref. Gralla et al. (2020). The parameters used are $(\rho,\Delta,\kappa)=(60,2.5,-2)$, with the first value corresponding to $\sim 6M$, the radial coordinate of the ISCO in Schwarzschild spacetime. Although the spacetime is non-rotating, the fluid velocity can be modeled with a mix of radial and (sub)Keplerian orbits, as described in e.g. Refs. Pu et al. (2016); Cárdenas-Avendaño et al. (2023). For the images in the main text, parameters $\xi=\beta_{r}=\beta_{\phi}=1.0$ in the conventions of Ref. Cárdenas-Avendaño et al. (2023) have been used, corresponding to pure Keplerian orbits which give rise to the observed asymmetry in the images.

### v2

@@ V Ray tracing
The ray-traced images presented in this work (cf.Fig. 6 of the main text) are obtained using Flexible Object-Oriented Ray Tracer (FOORT), which is publicly available on GitHub Mayerson ; Mayerson et al. . The geodesic equations are integrated using the Runge-Kutta 4 or the Verlet scheme to produce images of a variety of spacetimes. FOORT is highly flexible and easily accommodates new analytical or numerical spacetimes. It employs adaptive mesh refinement (AMR) by refining the initially uniform grid in regions where selected image features change significantly.
The images in the main text are produced for the S06A044 BS model using the Verlet scheme, for an observer located at areal radius $\mu r=1000$ in units where the mass of the boson star is $\mu M\approx 10.4$. The image region has dimensions $150\times 150$ in these units, and the highest resolution obtained with the AMR is $\sim 0.047$. The insets focus on a region of $7.5\times 7.5$, with highest resolution $\sim 0.0047$. The emission is modeled through an optically thin equatorial disk with emission profile given by Gralla et al. (2020)
$I_{0}(r)=\,$ $\left[(r-\rho)^{2}+\Delta^{2}\right]^{-1/2}$
@@
$\times\exp\left\{-\frac{1}{2}\left[\kappa+\operatorname{arcsinh}\left(\frac{r-\rho}{\Delta}\right)\right]^{2}\right\}\,.$ (S23)
@@
The parameters used are $(\rho,\Delta,\kappa)=(60,2.5,-2)$, with the first value corresponding to $\sim 6M$, the radial coordinate of the ISCO in Schwarzschild spacetime. Although the spacetime is non-rotating, the fluid velocity can be modeled with a mix of radial and (sub)Keplerian orbits, as described in Refs. Pu et al. (2016); Cárdenas-Avendaño et al. (2023). For the images in the main text, parameters $\xi=\beta_{r}=\beta_{\phi}=1.0$ (see Eqs. (B36), (B50), (B51) in Ref. Cárdenas-Avendaño et al. (2023)) have been used, corresponding to pure Keplerian orbits which give rise to the observed asymmetry in the images.